OpenStudy (anonymous):

MEDAL AND FAN FOR ANSWER What is the slope of the line passing through the points (–1, 3) and (4, –7)?

OpenStudy (anonymous):

do you have any answer choices

OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y? @cody7744 I need help with this one instead

OpenStudy (anonymous):

4x – 6y = 12 Divide by 2 just to make life simpler. 2x – 3y = 6 Subtract 2x –3y = -2x + 6 Divide by -3 y = (2/3)x - 2 Solving for y produces the slope-intercept form!! There's the slope you need! (0,-2) is the y- intercept. The slope is the other number.

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

so my answer is 2/3?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

do you need help with anything else